# RoundTable: Investigating Group Decision-Making Mechanism in Multi-Agent Collaboration

### v1

@@ 5 Results
5.1 Exchange Economy
The experiment results show distinct performance differences across social choice methods in Table 1 and Figure 2(a). Score-based mechanisms achieve higher performance early in the collaboration process compared to one-vote mechanisms, indicating that nuanced evaluations better aggregate agent preferences. One-vote mechanisms struggle with lower and less stable early performance due to increased disagreement, which impairs decision-making. Despite differing early trajectories, all social choice methods reach high-quality outcomes by the end. In most cases, performance peaks in the middle rounds, suggesting that early stopping could be advantageous. Rationality scores show that agents often make suboptimal choices, but decentralized MAS still performs well overall. Score-based mechanisms allow more dynamic decision updates, while one-vote mechanisms exhibit greater rigidity, particularly with unanimous voting.
When comparing one-vote mechanisms, unanimous voting performs the worst due to its inflexibility and need for total agreement, while majority voting slightly outperforms plurality. Additionally, fairness gradually decreases from unanimous to plurality voting. These observations suggest that, compared to the rigid unanimous method or the looser rules of plurality voting, a moderate level of decision flexibility leads to better outcomes, albeit with some sacrifice in fairness across agents.
For the results of ablation studies, please refer to Appendix D.1.
Table 1: Exchange economics environment results across different social choice methods. The reported numbers are an average of all simulations, and the numbers in parentheses are standard errors. The smallest and largest value in a category is colored in blue and red. We see that one-vote mechanisms show lower and less stable early performance, due to the higher probability of disagreement.
@@ 5.1 Exchange Economy
(a) Social choice comparison, using 3 agent, asymmetric, gpt-4o-mini setting.
(b) LLM comparison, using 3 agent, asymmetric, majority voting setting.
@@
Figure 2: Exchange economy environment results in group total utility. Line plots (left y-axis) show the group total utility achieved over rounds; bar plots (right y-axis) represent the ratio of cases where participants yet to reach an agreement until a certain round. The red horizontal line indicates the maximum achievable group total utility. We observe that decentralized collaboration performs well across various social choices and LLMs.
(a) Comparison between social choices in MAE.
(b) Comparison between social choices in RMSE.
@@
Figure 3: Comparison between social choices in recommendation system environment. Line plots (left y-axis) show the comparison of MAE/RMSE achieved over rounds; bar plots (right y-axis) represent the ratio of cases where participants yet to reach an agreement until a certain round.
5.2 Recommendation System
@@
We illustrate the performance of recommendation system in Figure 3(a), 3(b) and Table 7. The results shows that multi-agent collaboration achieves moderate performance with distributed information, with the decentralized MAS performing comparably to SoTA approaches. While the SoTA method was purely machine-learning-driven, the MAS approach showed reasonable outcomes, indicating that collaboration between agents can explore and produce meaningful signals that help to make decision. Multi-agent collaboration with access to SoTA models may lead to better performance, but we leave this experiment for future work. Similar to the exchange economy, multiple results in the recommendation system follow a V-shaped curve. This suggests that while agent collaboration was efficient early on, it declined in later rounds, highlighting diminishing returns from repeated interactions. The trend underscores the need to balance sustained collaboration within MAS over time.
@@
When comparing one-vote mechanisms to score-based mechanisms, we found that one-vote approaches generally worsened or plateaued after several rounds of collaboration. In contrast, score-based mechanisms showed consistent improvement over time, suggesting that score-based mechanisms are better suited for scenarios requiring continuous improvement, allowing for more nuanced input and refinement over multiple rounds.
5.3 Cross-Agent Conversation Analysis
5.3.1 Basic Statistics
Table 8 shows the average values of message length, complexity, and information difference per round. Message length increases over time in both environments, with shorter messages in early rounds and longer ones later. The recommendation system generally has longer messages due to its need for detailed explanations. Message complexity in the exchange economy starts low, peaks early, dips, and then gradually increases toward the final round, while the recommendation system shows consistently higher and increasing complexity throughout. Information difference steadily decreases in both environments, indicating a convergence of topics and less novel information as the discussions progress.
The dialogue act annotation results are presented in Table 9 in Appendix D.4. The dialogue act annotation results in both environments show that Inform and Request acts dominate, indicating frequent information sharing and input requests critical for task progression. The Confirm act rises sharply after round 3 in the recommendation system and more gradually in the exchange economy, showing validation is more prominent in recommendation tasks. Summarize acts increase slightly in later rounds, consolidating information, while Evaluate acts remain consistently high, reflecting ongoing assessment. Propose acts surge early in the exchange economy and peak later in the recommendation system, suggesting early proposals are vital in negotiations. Acts like Compromise, Defend, Accept, and Decline are rare, reflecting minimal adversarial behavior, while Others remain low, indicating the high quality of dialogue act definitions.
5.3.2 Dialogue Act Transition Graph
(a) Dialogue act transition graph for exchange economy.
(b) Dialogue act transition graph for recommendation system.
The most probable transition graphs in both environments are shown in Figure 4(a) and 4(b). The breakdown of social choices are displayed in Figure 5 and 6. The transitions reveal a structured progression from requesting information to proposing and evaluating solutions, followed by resolution or negotiation. In the exchange economy, collaboration centers around a loop between Request and Propose, reflecting the cooperative decision-making process. Key transitions include Inform to Request, Confirm to Request, and Compromise to Propose, showing how agents share knowledge, confirm information, and adopt ideas. Accept transitions to End, signaling negotiation closure. In contrast, the recommendation system centers on Request, indicating a more cooperative, exploratory process with incomplete information.
5.4 Early Stopping in MAS
Table 2: Comparison between early stopping methods in recommendation system environment. The performance is compared with Oracle and baseline @10 in MAE. ($\downarrow$) indicates better performance with higher and lower values, respectively. Experiments are based on the results of 3 agents, gpt-4o-mini setting. Results are based on 5-fold cross validation. We observe that language-based methods performed well overall.
@@ 5.4 Early Stopping in MAS
We show comparison of early stopping methods in recommendation system in Table 2, full results in Table 10, and their basic statistics are in Table 11 in Appendix D.5. To begin, V-shaped performance trends were observed across most cases, where the best outcome is observed before reaching the final checkpoint (@10), indicating that intermediate stopping leads to better results than allowing the process to fully complete. This pattern supports the necessity of early stopping in multi-agent collaboration settings, where overextending the interaction can lead to diminishing returns in performance.
Almost all early stopping methods outperformed the baseline (@10), reinforcing the effectiveness of early termination in improving outcomes. Notably, methods leveraging linguistic features, such as Information Difference and Dialogue Act, delivered better performance across different social choices and environments. These findings highlight the advantage of incorporating dialogue-based metrics in deciding when to stop, especially in complex multi-agent environments. Their results across both the exchange economy and recommendation system scenarios further supports the claim that linguistic indicators can be powerful tools for optimizing collaboration outcomes, reducing cognitive load, and improving decision efficiency.
@@ 6 Discussion
6.1 Impact of Social Choice Methods on Collaborative Behavior
@@
The study demonstrates that social choice methods significantly impact multi-agent collaboration, with score-based mechanisms achieving higher performance and efficiency, especially in early rounds, by allowing agents to express nuanced preferences. In contrast, one-vote mechanisms struggle with rigidity, leading to lower initial performance. The strictness of social choice methods, such as unanimous voting, reduces collaboration performance due to its inability to accommodate dissent. Score-based mechanisms, which support expressing preference gradients, foster dynamic exchanges and improved decision quality, making them more effective in negotiation-heavy tasks.
@@
One-vote mechanisms also have unique advantages. First, the format of voting is simpler than score-based mechanisms, which led to fewer format errors during experiments from LLMs. Furthermore, unanimous voting showed the highest fairness and non-decreasing outcome over rounds when utility is quantifiable. These strengths suggest that social choice methods in decentralized MAS should be thoughtfully chosen based on the specific task, or the primary metric being prioritized. Mixture of different methods in different stages of collaboration may also bring effectiveness, which we leave for future studies.
6.2 Linguistic Indicators of Collaboration
@@
Linguistic analysis of agent conversations reveals several key indicators of effective collaboration. The length and complexity of messages increase progressively across rounds, suggesting that deeper engagement and richer exchanges occur as the collaboration advanced. This is particularly evident in the recommendation system environment, where tasks require greater information exchange and negotiation. Complex tasks inherently demand more intricate language, as agents need to convey not just factual information but also interpret, analyze, and synthesize inputs from others.
@@
Dialogue acts provide another layer of insight into collaborative behavior. Inform, Propose and Request acts dominate conversations, highlighting that agents were frequently sharing new information and seeking clarifications, which are essential for progressing towards a solution. Confirm act became more prevalent in later stages, particularly in environments where validation was crucial. Interestingly, Propose and Evaluate acts were more frequent in negotiation-oriented tasks, like exchange economies, where agents needed to offer and assess potential solutions regularly. This linguistic pattern underscores that successful collaboration hinges on the flow of information, clarification, and ongoing evaluation of proposals. These indicators align with patterns seen in human collaboration, where information exchange and continuous feedback loops drive cooperative success.
6.3 Early Stopping in Multi-Agent Collaboration
@@
Early stopping is essential in multi-agent collaboration because it prevents diminishing returns and inefficiency in decision-making. Our experiments shows that while the initial rounds of collaboration led to significant improvements in group utility and decision quality, continuing beyond a certain point often results in stagnant or even declining performance. This highlights the need for mechanisms that can intelligently terminate the collaboration process once optimal solutions have been reached, avoiding unnecessary iterations.
@@
Key insights into effective early stopping methods emerge from analyzing conversational patterns. Information difference suggests that as the rounds progress, the novelty of information decreases, with fewer new ideas being introduced. Transition graphs of dialogue acts reveal common collaborative patterns. These signals lead to linguistic feature-based stopping methods, which halt collaboration when no meaningful new information is detected, or when dialogue act patterns indicate the need for termination. Our findings demonstrate that implementing early stopping methods based on linguistic cues can enhance the efficiency and effectiveness of decentralized multi-agent collaboration by preventing unnecessary prolonged discussions.
@@ 7 Conclusion
This paper investigates the dynamics of cross-agent communication and decentralized decision-making in MAS, exploring how agent conversation and social choice methods impact collaboration and collective intelligence. We show that only the correct selection of a social choice method for a given environment can promote the success of a MAS, linguistic features of agent conversations serve as indicators of effective collaboration, and these cues provide valuable information for the effective termination of MAS.
LLM generation and prediction are unstable, with single agents often making quasi-random actions based on observations. Decentralized MAS improve performance by guiding agents through conversations to make more targeted proposals and enabling group decision-making to select the best option. Our study provides insights into how group communication and decentralized decision-making enhance multi-agent collaboration, offering important implications for designing more efficient MAS environments.
Acknowledgments
We would like to thank Dr. Lyle Ungar, Dr. Sharath Guntuku, and Dr. Dan Roth for their thoughtful advice. This work was partly supported by a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI.
Appendix A Details of RoundTable
A.1 RoundTable Platform Design
@@
Input and Initialization defines each agent for collaboration. Each agent will be given a specific $C_{i}$ and $u_{i}$, which include the same input user query, information of RoundTable, and the social choice method $F$. However, they may have different background, information access, or utility function. While $C_{i}$ and $u_{i}$ can be defined automatically by agent recruiting or profiling (Chen et al., 2023), we manually define them for simplicity.
Message Phase is the initial stage of the iterative collaboration process, where all agents participate in a simultaneous, open chat conversation. During this phase, agents can freely choose recipients based on their context to send messages. They can respond to others’ questions, share insights, or ask their own questions. There are no restrictions on whom to communicate with or what to say. Once generated, all messages are updated to $C_{i}$ for everyone, regardless of the original sender or intended recipients. This process guarantees that all messages are sent simultaneously, makes it has no specific order of communication. Since only one message can be sent per iteration, back-and-forth conversations can take place across multiple rounds.
Proposal Phase allows agents to present a proposal $p_{i}$, offering a potential solution or response to the user query from their perspective, aiming to fulfil their own goal or maximize the utility. Before generating a proposal, agents will be asked to do a step by step reasoning, which is by default invisible to other agents. Reasoning is designed to give agents a chance to summarize and analyze information in $C_{i}$. The most recent proposal from each agent will be considered as a candidate for voting in the next phase. Like the Message Phase, proposals are made simultaneously and will be updated for $C_{i}$ once everyone has generated a response. Agents also have the option to skip submitting a new proposal; in such cases, their previous proposal will automatically be included among the candidates.
Voting Phase collects one latest proposal $p_{i}$ from each agent, additionally include the latest intermediate group decision $x^{*}$ from previous rounds to form a candidate proposal list. Agents in this round are asked to express preferences toward candidates. The preferences can be in various formats, including voting, rating and ranking, depends on what types of social choice method $F$ is used. Similar with the proposal phase, agents have option to not vote (give up). With all the collected proposals and preferences, social choice method selects one proposal as the intermediate group decision. Finally, voting details and the result are updated to all agents’ context $C_{i}$.
@@
A.2 RoundTable Prompt Design
We present the prompt design for RoundTable in this section. Italicized words represent inputs or variables within the prompt.

### v2

@@ 5 Results
Table 1: Exchange economics environment results across different voting rules. The reported numbers are an average of all simulations, and the numbers in parentheses are standard errors. The smallest and largest value in a category is colored in blue and red. We see that discrete voting rules show lower and less stable early performance, due to the higher probability of disagreement.
Exchange Economy
The experiment results show distinct performance differences across voting rules in Table 1 and Figure 3.777We explore more details with ablation studies, including scalability test with different number of agents and generalizability experiment with varying LLM models. For the results, please refer to Appendix E.1. Preference voting rules achieve higher performance early in the collaboration process compared to discrete voting rules, indicating that nuanced evaluations better aggregate agent preferences. Discrete voting rules struggle with lower and less stable early performance due to greater disagreement, which impairs decision-making. Despite differing early trajectories, all voting rules reach high-quality outcomes by the end. In most cases, performance peaks in the middle rounds, suggesting that early stopping could be advantageous. Rationality scores show that agents often make suboptimal choices, but decentralized MAS still performs well overall. Preference voting rules allow more dynamic decision updates, while discrete voting rules exhibit greater rigidity, particularly with unanimous voting.
When comparing discrete voting rules, unanimous voting performs the worst due to its rigidity and requirement for total agreement, resulting in an 87% lower initial round performance and a 39% lower final round performance compared to the best-performing models. Moreover, fairness progressively declines from unanimous to plurality voting. These findings suggest that stricter voting thresholds, such as unanimity, enforce greater fairness among agents but lead to significantly worse outcomes.
@@
Figure 3: Voting rules comparison in exchange economy, using 3 agent, asymmetric, gpt-4o-mini setting. Line plots (left y-axis) show the group total utility achieved over rounds; bar plots (right y-axis) represent the ratio of cases where participants yet to reach an agreement until a certain round. The red horizontal line indicates the maximum achievable group total utility.
@@
Figure 4: Comparison between voting rules in recommendation system environment. Line plots (left y-axis) show the comparison of MAE achieved over rounds; bar plots (right y-axis) represent the ratio of cases where participants yet to reach an agreement until a certain round. The green horizontal line is baseline (Always Guess 4), and the blue line is SoTA performance Behera and Nain (2023).
Recommendation System
@@
We illustrate the performance of recommendation system in Figure 4, 21 and Table 7. The results show that multi-agent collaboration achieves moderate performance with distributed information, with the decentralized MAS performing comparably to SoTA approaches. While the SoTA method was purely machine-learning-driven, the MAS approach showed reasonable outcomes, indicating that collaboration between agents can explore and produce meaningful signals that help to make decision. Similar to the exchange economy, multiple results in the recommendation system follow a V-shaped curve. This suggests that while agent collaboration was efficient early on, it declined in later rounds, highlighting diminishing returns from repeated interactions. The trend underscores the need to balance sustained collaboration within MAS over time.
@@
When comparing discrete to preference voting rules, we find that discrete voting rules generally worsened or plateaued after several rounds of collaboration. In contrast, preference voting rules show consistent improvement over time. This suggests that adaptive and flexible voting formats enhance collaboration by allowing for more nuanced input and iterative refinement across multiple rounds.
6 Cross-Agent Communication Analysis
Method
From the experiment results, we notice the performance degraded throughout the prolonged collaboration. To understand agent collaboration behaviors and communication, we perform a language analysis focused on three key features. [^8] Detailed definition of four features is shown in Appendix D.4. All observations are collected in 3 agents, gpt-4o-mini setting. Message Length: The length of the message. Message Complexity: The complexity of the message. Information Distance: The embedding distance between the conversations from current and previous round. Dialogue Acts: The actions performed in the conversation, annotated with LLM. We extracted features from each round of communication and analyzed the trend of changes across different rounds, voting rules, and environments.
Figure 5: Line plots of average message length and information distance per round. Increasing message length and decreasing information distance indicate collaboration inefficiency in prolonged rounds.
Results
Table 8 and Figure 5 shows the averages of language features per round. There are two notable points: 1) Recommendation system is a harder environment than exchange economy, presented from overall higher message length and complexity. 2) Message length increases by 84% over time, while information distance steadily decreases to 10% in both environments. These phenomena indicate a convergence of topics and a decline in novel information as discussions progress, while the increasing message length adds additional cognitive and communication costs.
Dialogue acts annotation results are presented in Table 9, and the transition graph between dialogue acts is in Figure 22(a), 22(b), 23 and 24.999Detailed explanation of result can be found in Appendix E.3. In both environments, information exchange (Inform and Request) is the most frequent action, while direct disagreement (Defend and Decline) is very rare. Increasing usage of Confirm, Summarize and Compromise also indicates the progress of collaboration. Dialogue acts also show differences across environments. Exchange economy shows more Evaluate, Propose, and strong loop between Request and Propose, showing how agents collaborate when the relationship is competitive. Meanwhile, recommendation system shows minimal compromise and accept, which are unnecessary in cooperate relationship. Instead, it has intensive use of request, indicating the focus on data exploration.
7 Early Stopping in MAS
UNA. MAJ. PLU. RAT. RAN. CUM.
@@ 7 Early Stopping in MAS
Table 2: Comparison between early stopping methods in recommendation system environment. The performance is compared with Oracle and baseline @10 in MAE. ($\downarrow$) indicates better performance with lower values. Experiments are based on the results of 3 agents, gpt-4o-mini setting. Results are based on 5-fold cross validation. We observe that language-based methods performed well overall.
Method
While enforced 10 rounds of collaboration to study agent behaviors, we observed signals of ineffective collaboration from task performance and language feature analysis. To enhance collaboration efficiency, we compare five early stopping methods: First Agreement, Consecutive Agreements, Validation Checkpoint, Information Distance, and Dialogue Act. [^10] Detailed definitions of early stopping methods are listed in Appendix D.7 We focus on the performance of Information Distance and Dialogue Act to examine whether language features help in identifying early stopping signals.
Results
We show comparison of early stopping methods in recommendation system in Table 2, full results in Table 10, and their basic statistics are in Table 11.
Almost all early stopping methods outperformed the non-stop baseline (@10), indicating effectiveness of early termination in improving outcomes. Notably, methods leveraging language features (Information Distance and Dialogue Act) delivered better performance across most of voting rules and environments, each closed 12% and 15% of the gap between the non-stop and oracle approaches with fewer rounds. These findings highlight the advantage of incorporating language-based metrics in deciding when to stop, especially in complex multi-agent environments. Their results across both the exchange economy and recommendation system environments further supports the claim that language indicators can be powerful tools for optimizing collaboration outcomes, reducing cognitive load, and improving decision efficiency.
@@ 8 Discussion
Impact of Voting Rules on Collaborative Behavior
@@
This study demonstrates that voting rules, particularly their thresholds and formats, play a crucial role in multi-agent collaboration. Preference voting rules with a flexible format achieve higher performance and efficiency, especially in early rounds. In contrast, discrete voting rules struggle with rigidity, leading to lower initial performance. The strictness of voting thresholds, such as in unanimous voting, hampers collaboration due to its inability to accommodate dissent. Overall, preference voting rules, which allow for nuanced preference expression, facilitate dynamic exchanges and enhance decision quality, making a collaboration more effective for complex tasks.
@@
Discrete voting rules also have unique advantages. First, the format of voting is simpler than preference voting rules, which led to fewer format errors during experiments from LLMs. Furthermore, unanimous voting showed the highest fairness and non-decreasing outcome over rounds when utility is quantifiable. These strengths suggest that voting rules in decentralized MAS should be thoughtfully chosen based on the specific task, or the primary metric being prioritized. Mixture of different methods in different stages of collaboration may also bring effectiveness, which we leave for future studies.
Language Indicators of Collaboration
@@
Language analysis of agent conversations reveals several key indicators of effective collaboration. The length and complexity of messages increase progressively across rounds, suggesting that deeper engagement and richer exchanges occur as the collaboration advanced. This is particularly evident in the recommendation system environment, where tasks require greater information exchange and negotiation. Complex tasks inherently demand more intricate language, as agents need to convey not just factual information but also interpret, analyze, and synthesize inputs from others. A key takeaway is the decline in collaboration efficiency over time, indicated by increasing message length and decreasing information distance between adjacent rounds.
@@
Dialogue acts provide another layer of insight into collaborative behavior. We discover that dialogue acts comprehensively reflects agents’ collaboration behaviors. When competitive relationship makes agents do more negotiation, cooperate relationship stimulate information exchange. This language pattern underscores that successful collaboration hinges on the flow of information, clarification, and ongoing evaluation of proposals. These indicators align with patterns seen in human collaboration, where information exchange and continuous feedback loops drive cooperative success.
Early Stopping
@@
As a part of decision-making mechanism, early stopping is essential in multi-agent collaboration because it prevents diminishing returns and inefficiency in decision-making. Our experiments show that while the initial rounds of collaboration led to significant improvements in group utility and decision quality, continuing beyond a certain point often results in stagnant or even declining performance. This highlights the need for mechanisms that can intelligently terminate the collaboration process once optimal solutions have been reached, avoiding unnecessary iterations.
@@
Key insights into effective early stopping methods emerge from analyzing conversational patterns. Information distance suggests that as the rounds progress, the novelty of information decreases, with fewer new ideas being introduced. Transition graphs of dialogue acts reveal common collaborative patterns. These signals lead to language feature-based stopping methods, which halt collaboration when no meaningful new information is detected, or when dialogue act patterns indicate the need for termination. Our findings demonstrate that implementing early stopping methods based on language cues can enhance the efficiency and effectiveness of decentralized multi-agent collaboration by preventing unnecessary prolonged discussions.
@@ 9 Conclusion
This paper investigates the role of group decision-making mechanisms in decentralized MAS, emphasizing how different voting rules shape collaboration dynamics and efficiency. Through controlled experiments on RoundTable, we analyze the impact of decision thresholds and formats on performance, finding that adaptive voting rules lead to more effective outcomes. Our analysis of agent communication patterns reveals increasing message length and repetition over time, indicating inefficiencies in prolonged collaboration. Building on these insights, we demonstrate that language-based early stopping improves performance while reducing unnecessary rounds. Our study provides insights into decision-making and cross-agent communication, offering valuable implications for improving multi-agent collaboration and efficiency.
Limitations
Platform Design
In RoundTable, we fixed the phase order in our platform design to follow a structured sequence: message, propose, and voting. While this sequence represents a common collaboration flow, alternative action chains are possible, such as multi-round discussions before voting or iterative proposal refinements. However, we designed RoundTable to ensure that it captures the essential decision-making steps while maintaining clarity and reproducibility for experimental evaluation. By including both proposal and voting phases in each round, our design not only provides better transparency of agent intent but also allows researchers to collect granular collaboration data at multiple stages. This structured approach facilitates analysis of how agents form consensus and adapt strategies over time. Furthermore, repeating round-based collaboration will eventually enable agents to engage in multiple rounds of communication, proposing, and voting, giving them the chance to continue their discussion based on previous interactions. Although other collaboration models exist, we believe RoundTable represents a generalizable framework for studying the impact of decision-making mechanisms in decentralized MAS.
Environment Selection
We conducted our experiments in two distinct environments: exchange economy and recommendation system. These environments were chosen because they inherently involve a balance between competition and cooperation, making them well-suited for studying collaborative decision-making. In exchange economy, agents must negotiate and trade resources while optimizing individual and collective goals, capturing the tension between self-interest and group benefit. In contrast, recommendation system simulates a setting where agents collaborate to improve recommendations based on user preferences, emphasizing mutual benefit and shared optimization. While our study focuses on these two environments, they represent a broad spectrum of real-world multi-agent interactions, spanning both competitive resource allocation and collaborative preference aggregation scenarios. As such, our findings offer insights applicable to a wide range of collaborative decision-making contexts.
Appendix A Details on the Background of Decentralized MAS
Agents powered by LLMs have demonstrated impressive problem-solving capabilities across a wide range of tasks. However, single-agent systems struggle with problems that are too large or complex, often leading to instability, misalignment with the intended request, and hallucination (Liu et al., 2024; Kuhn et al., 2023; Lyu et al., 2023). To overcome these limitations, research has increasingly turned to Multi-Agent Systems (MAS). MAS leverage collective intelligence by enabling agents to specialize in distinct skills and collaborate effectively (Guo et al., 2024).
In MAS, agents often hold differing perspectives due to variations in persona roles, information access, and individual goals, making group decision-making inherently complex. Many LLM-driven MAS rely on centralized decision-making, where a manager agent or predefined power hierarchy resolves conflicts (Hong et al., 2023; Qian et al., 2023; Dong et al., 2023). However, centralization introduces several challenges: 1) fairness: Individual agent inputs may be misrepresented in the final decision (Jiang and Lu, 2019; Owens et al., 2024); 2) rigidity: Fixed hierarchical structures limit adaptability, making systems prone to overfitting (Chen et al., 2023); 3) information constraints: A central authority cannot always dictate optimal strategies when agents possess private or partial information (Xu et al., 2023).
Decentralized group decision-making mitigates the limitations of centralized MAS by distributing decision power among agents, allowing each to actively participate in the process. This approach is particularly critical in dynamic and complex environments, where agents must operate independently with information asymmetry (Mandi et al., 2024; Zhang et al., 2023; Xu et al., 2023; Park et al., 2023). By removing reliance on a central authority, decentralized MAS enable greater adaptability and resilience. However, in the absence of a fixed decision-maker, agents must establish structured yet flexible coordination strategies to achieve coherent and effective group decisions. Identifying mechanisms that facilitate meaningful agreement in such systems remains a key challenge.
Appendix B Details of RoundTable
B.1 RoundTable Platform Design
@@
Input and Initialization defines each agent for collaboration. Each agent will be given a specific $\Omega_{i}$ and $u_{i}$, which include the same input user query, information of RoundTable, and the voting rule $F$. However, they may have different background, information access, or utility function. While $\Omega_{i}$ and $u_{i}$ can be defined automatically by agent recruiting or profiling (Chen et al., 2023), we manually define them for simplicity.
Message Phase is the initial stage of the iterative collaboration process, where all agents participate in a simultaneous, open chat conversation. During this phase, agents can freely choose recipients based on their context to send messages. They can respond to others’ questions, share insights, or ask their own questions. There are no restrictions on whom to communicate with or what to say. Once generated, all messages are updated to $\Omega_{i}$ for everyone, regardless of the original sender or intended recipients. This process guarantees that all messages are sent simultaneously, makes it has no specific order of communication. Since only one message can be sent per iteration, back-and-forth conversations can take place across multiple rounds.
Proposal Phase allows agents to present a proposal $p_{i}$, offering a potential solution or response to the user query from their perspective, aiming to fulfil their individual objective or maximize the utility. Before generating a proposal, agents will be asked to do a step by step reasoning, which is by default invisible to other agents. Reasoning is designed to give agents a chance to summarize and analyze information in $\Omega_{i}$. The most recent proposal from each agent will be considered as a candidate for voting in the next phase. Like the Message Phase, proposals are made simultaneously and will be updated for $\Omega_{i}$ once everyone has generated a response. Agents also have the option to skip submitting a new proposal; in such cases, their previous proposal will automatically be included among the candidates.
Voting Phase collects one latest proposal $p_{i}$ from each agent, additionally include the latest intermediate group decision $x^{*}$ from previous rounds to form a candidate proposal list. Agents in this round are asked to express preferences toward candidates. The preferences can be in various formats, including voting, rating and ranking, depends on what types of voting rule $F$ is used. Similar with the proposal phase, agents have option to not vote (give up). With all the collected proposals and preferences, voting rule selects one proposal as the intermediate group decision. Finally, voting details and the result are updated to all agents’ context $\Omega_{i}$.
@@
B.2 RoundTable Prompt Design
We present the prompt design for RoundTable in this section, which consists of initialization stage (Figure 6), message phase (Figure 7), proposal phase (Figure 8), and voting phase (Figure 9). Italicized words represent inputs or variables within the prompt.